solve each system using substitution.

2=2y-x
23=5y-4x

Answer by mananth(16946)

2=2y-x
x=2y-2
23=5y-4x
substitute x in second equation
23=5y-4x
23=5y-4(2y-2)
23=5y-8y+8
23=-3y+8
-3y=23-8
y=-5
plug y in equation to get x
x=2y-2
x=-10-2
x=-12